### v2

# VPO: Aligning Text-to-Video Generation Models with Prompt Optimization

Abstract
Video generation models have achieved remarkable progress in text-to-video tasks. These models are typically trained on text-video pairs with highly detailed and carefully crafted descriptions, while real-world user inputs during inference are often concise, vague, or poorly structured. This gap makes prompt optimization crucial for generating high-quality videos. Current methods often rely on large language models (LLMs) to refine prompts through in-context learning, but suffer from several limitations: they may distort user intent, omit critical details, or introduce safety risks. Moreover, they optimize prompts without considering the impact on the final video quality, which can lead to suboptimal results. To address these issues, we introduce VPO, a principled framework that optimizes prompts based on three core principles: harmlessness, accuracy, and helpfulness. The generated prompts faithfully preserve user intents and, more importantly, enhance the safety and quality of generated videos. To achieve this, VPO employs a two-stage optimization approach. First, we construct and refine a supervised fine-tuning (SFT) dataset based on principles of safety and alignment. Second, we introduce both text-level and video-level feedback to further optimize the SFT model with preference learning. Our extensive experiments demonstrate that VPO significantly improves safety, alignment, and video quality compared to baseline methods. Moreover, VPO shows strong generalization across video generation models. Furthermore, we demonstrate that VPO could outperform and be combined with RLHF methods on video generation models, underscoring the effectiveness of VPO in aligning video generation models. Our code and data are publicly available at https://github.com/thu-coai/VPO.
Figure 1: (Upper) The critical gap between the training and inference stages in video generation models. During training, video descriptions are detailed, descriptive, and well-crafted, while user inputs in the inference stage are often concise, unclear, and poorly structured, creating a mismatch that hampers model performance. (Lower) Comparison of VPO with traditional prompt optimization methods. Traditional methods rely on the in-context learning capabilities of LLMs, which can lead to issues such as safety issues, misalignment, and low-quality videos. In contrast, VPO serves as a harmless, accurate, and helpful prompt optimizer for high-quality video generation.
1 Introduction
Recent advancements in text-to-video generation models have significantly improved the ability to produce high-quality video content [4, 18, 15, 29, 28]. These models typically rely on extensive, high-quality video-text annotated datasets, where detailed labels capture specific video content. While training descriptions are detailed and carefully crafted, real-world user inputs are often brief, ambiguous, or poorly structured. This critical gap between model training and inference can lead to suboptimal outputs, making it crucial to optimize and refine user inputs to generate desired high-quality video outputs (Figure 1).
Current approaches to optimizing user inputs for video generation mostly utilize large language models (LLMs) with in-context learning [28, 29, 15]. However, these methods suffer from several limitations:
  • Safety concerns: Existing methods do not explicitly ensure that optimized prompts remain safe, potentially leading to inappropriate or harmful video content.
  • Imprecise refinement: Current methods may unintentionally alter the user’s intent or introduce biases, leading to outputs that deviate from the original query.
  • Neglecting final video quality: Current approaches optimize prompts to be semantically richer, but do not explicitly consider how the refined prompt impacts the actual generated video, limiting their practical effectiveness.
To address these challenges, we propose a set of alignment principles for video prompt optimization, inspired by the HHH (Harmless, Honest, Helpful) principle in LLMs [1]. An aligned prompt optimization model should adhere to the following:
  • Harmless: The refined prompt should avoid any harmful content, such as depictions of blood or violence.
  • Accurate: The optimized prompt should be precisely aligned with the user input, except for safety issues.
  • Helpful: The refined prompt should be detailed and descriptive, and can help the video generation model produce high-quality video content.
Building on these principles, we introduce VPO, a principle-driven framework for optimizing user inputs for text-to-video generation, which integrates two critical stages: principle-based supervised fine-tuning (SFT) and multi-feedback preference optimization. In the SFT stage, we first leverage the in-context learning capabilities of LLMs to construct an initial SFT dataset. This data is then improved through the application of principle-based critique and refinement, ensuring the generated prompts are safe, accurately aligned with the original query, and sufficiently detailed. The resulting fine-tuned model poses a foundational ability to generate prompts following the proposed principles. In the subsequent preference optimization phase, we incorporate both text-level feedback and video-level reward signals to further improve the SFT model. By combining these feedback sources, we construct preference data that prioritizes harmlessness, accuracy, and helpfulness in the optimized prompt. Using Direct Preference Optimization (DPO), we train the model to generate prompts that consistently result in safer, more precise, and higher-quality video outputs. This two-stage framework explicitly integrates alignment principles into the prompt optimization process, bridging the critical gap between model training and real-world inference in video generation.
We have conducted experiments on several popular video generation models, including CogVideoX [28] and Open-Sora [29]. Through extensive experiments, we demonstrate the effectiveness of VPO in generating harmless, accurate, and helpful prompts for high-quality video content, outperforming traditional prompt rewriting methods. Notably, on CogVideoX, VPO improves the win rate of 37.5% over original user queries and 14% over the official prompt optimization method in human evaluation. Moreover, VPO significantly reduces the unsafety rate, enhancing the overall harmlessness of video generation models. We also highlight the necessity of multi-feedback preference optimization, demonstrating its crucial role in refining prompts for improved video generation. Furthermore, we demonstrate the potential of VPO as an RLHF method. VPO not only surpasses Diffusion DPO [25], but also brings additional benefits over Diffusion DPO. Additionally, VPO exhibits good generalization capabilities across different models—when trained on CogVideoX, it enhances the performance of Open-Sora 1.2, indicating the potential for training a general prompt optimizer for diverse video generation models.
Our contributions can be summarized as follows:
  • We propose the principle of the alignment of video prompt optimizer: harmless, accurate, and helpful. Based on this, we introduce VPO, a systematic framework that integrates principle-based supervised fine-tuning and multi-feedback preference optimization, to build an aligned video prompt optimizer.
  • We demonstrate the effectiveness of VPO for improving query refinement accuracy, enhancing the safety of generated videos, and boosting the overall performance of video generation models. Furthermore, VPO exhibits good generalization across different video generation models.
  • We demonstrate that VPO is a novel and competitive alignment method for text-to-video models, outperforming Diffusion DPO and providing orthogonal benefits.
2 Related Work
2.1 Video Generation
With the widespread application of diffusion models in the domain of image generation [21, 20], research efforts have increasingly shifted towards video generation models [3, 8, 23]. Video Diffusion Models [11] firstly extend image diffusion models to video generation by employing a 3D UNet architecture. Text2Video-Zero [14] involves motion dynamics in image generation models to enable zero-shot video generation. However, these approaches encounter significant challenges in generating long videos. Subsequent studies, such as Stable Video Diffusion [2] and VideoCrafter [5], leverage large-scale pre-training on high-quality video datasets, yet their scalability remains constrained due to the limitation of architectures like UNet. By adopting Transformers [24] as the backbone of diffusion models, the scalable architecture of DiT [17] has substantially enhanced the capability of video generation. Building on DiT, several impressive video generation models have emerged [4, 28, 18, 15]. Nonetheless, the training data for these video generation models typically consists of detailed video descriptions, whereas user inputs during inference are often brief, unstructured, or even ambiguous in intent. Bridging this gap is critical for improving the quality of generated videos, underscoring the importance of training a robust prompt optimization model.
2.2 Prompt Optimization
Prompt optimization techniques have long been a pivotal research problem for generation models. Early work on automatic prompt optimization can be traced back to AutoPrompt [22], which utilizes gradient-guided search to automatically create prompt templates. With the rapid development of LLMs, there has been a significant increase in studies leveraging LLMs for automatic prompt optimization [30, 6]. Similarly, existing video prompt optimization methods employ LLMs to refine user inputs, typically through in-context learning [28, 15, 29]. However, these methods rely solely on the inherent capabilities of LLMs, without considering the alignment with user query and the quality of generated videos, thus often resulting in suboptimal video outputs. Prompt-A-Video [13] incorporates image and video reward models during the optimization process to enhance the quality of generated videos, but it fails to account for text-level alignment, which is essential for ensuring the safety and accuracy of rewritten prompts. To address these limitations, we propose VPO, a systematic framework for constructing a video prompt optimization model that can safely and accurately capture user intent, and assist video generation models in generating high-quality videos.
3 Method
Figure 2: The overall framework of VPO. VPO consists of two main steps: Principle-Based SFT and Multi-Feedback Preference Optimization. By refining user queries into harmless, accurate, and helpful prompts, VPO acts as a bridge between users and video generation models, ultimately enhancing the quality of generated videos.
We introduce VPO, an automated framework designed to bridge the gap between real-world user inputs and the well-crafted text descriptions used during model training. Unlike existing approaches that rely solely on the in-context learning capabilities of LLMs, VPO follows a principle-driven approach, ensuring that prompt refinements are harmless, accurate, and helpful. By incorporating these core principles into both supervised fine-tuning and preference optimization, VPO effectively transforms concise, unclear, or unsafe user inputs into detailed, well-structured, and safe prompts that significantly enhance video generation quality.
3.1 Overall Framework
The overall framework of VPO is illustrated in Figure 2. Our framework consists of two key stages: Principle-Based SFT (§3.2) and Multi-Feedback Preference Optimization (§3.3). In the first stage, Principle-Based SFT, we construct a high-quality dataset to train an initial model that refines user inputs while adhering to our three core principles. This stage provides the model with the foundational capabilities to optimize user input into a harmless, accurate, and helpful prompt for text-to-video generation. In the second stage, Multi-Feedback Preference Optimization, we further improve the SFT model using both text-level and video-level feedback. Text-level feedback helps refine prompts to be well-aligned with user intent while maintaining harmless. Meanwhile, video-level feedback ensures that the generated prompts can lead to high-quality video outputs, effectively enhancing the performance of the text-to-video model.
3.2 Principle-Based SFT
We first construct a high-quality SFT dataset comprising user query and optimized prompt pairs to initialize the prompt optimizer with the ability to generate prompts that align with our guiding principles.
Query Curation.
To build the SFT dataset, we begin by collecting a comprehensive set of user queries. We utilize the VidProM dataset [26], which contains over one million real-world text-to-video queries. We adopt its deduplicated version and apply additional filtering steps to enhance the data. First, we perform rule-based quality filtering, considering factors such as keywords, special characters, and query length. To ensure query diversity, we employ the self-BLEU [31] metric to filter out overly similar queries. Additionally, since we aim to develop a prompt optimizer that prioritizes safety, we incorporate safety-related queries into our dataset. Specifically, we extract prompts labeled as unsafe in the original dataset. However, as Detoxify [9] introduces labeling errors, we re-evaluate these safety-related queries using a more robust LLM. Following this data curation process, we finalize a dataset consisting of approximately 18k general queries and 2k safety-related queries for training, where 10k are used for SFT and 10k for DPO.
Optimized Prompts Construction.
With a diverse and high-quality set of user queries, we employ an LLM, specifically GPT-4o, with in-context learning to generate initial optimized prompts. The demonstrations used in this process are carefully crafted to guide the LLM in producing clear, well-structured prompts. Additionally, we instruct the LLM to account for potential safety concerns, ensuring that the generated prompts avoid harmful content, such as depictions of violence or explicit material. The exact prompt template used in this data construction process is provided in the Appendix.
Principle-Based Refinement.
After generating query-prompt pairs $(x,p)$, where $x$ is the user query and $p$ is the optimized prompt, we further refine the prompts using a principle-driven approach to enhance their quality. We employ an LLM-as-a-judge method to assess prompts based on three key principles: harmlessness, accuracy, and helpfulness. The LLM critiques the generated prompts, identifying issues such as harmful content, missing key details from the user query, or vague scenario descriptions. We then collect these critiques and represent problematic cases as triplets $(x,p,c)$, where $c$ is the critique. Based on this critique, we refine the generated prompt $p$ to produce a refined version $p_{\text{refined}}$.
Model Training.
With the principle-based prompt refinement, we construct a high-quality SFT dataset, denoted as $(x,s)\in D_{\text{SFT}}$. For the prompt without being pointed out issues within critiques, $s=p$, retaining the LLM-generated prompt. If issues are identified, we use the refined version, setting $s=p_{\text{refined}}$. To initialize the prompt optimizer, we apply standard supervised fine-tuning with the loss function:
$\mathcal{L}=-\frac{1}{N}\sum_{i=1}^{N}\text{log}P(s|x,s_{<i}),$ (1)
where $N$ represents the length of $s$.
3.3 Multi-Feedback Preference Optimization
After obtaining the SFT model, we further enhance its capability in optimizing text-to-video generation prompts. Following our proposed principles, we incorporate two types of feedback. Text-level feedback assesses whether prompts align with user intent and maintain safety, while video-level feedback ensures that prompts result in high-quality video generation. Combining feedback from both dimensions, we can perform preference optimization on the model, allowing it to generate prompts that are safer, more accurate, and effective.
Data Sampling.
To support the DPO training, for each input $x$, we first sample $K$ optimized prompts $(p_{1},p_{2},\dots,p_{K})$ from the SFT model. We then construct DPO training pairs based on both text-level and video-level feedback.
Text-Level Preference Data Construction.
A desired prompt optimizer should produce prompts that are harmless, faithfully represent the user’s intent, and are clear and descriptive. To achieve this, in the preference optimization stage, we leverage an LLM-as-a-judge approach to provide text-level feedback. Given a user query and an optimized prompt pair $(x,p_{j})$, the LLM generates a critique to check whether the prompt violates any proposed principles. If a flaw is detected, we refine the prompt based on the critique, producing a revised version $p_{j_{\text{refined}}}$. These pairs $(x,p_{j}<p_{j_{\text{refined}}})$ comprise $D_{\text{text}}$.
Video-Level Preference Data Construction.
Ultimately, the helpfulness of a prompt should be judged by the quality of the video it helps generate. To guide the prompt optimizer in producing high-quality outputs, we utilize VisionReward [27], a state-of-the-art video reward model, to provide automated feedback. For prompts that pass the text-level check, we generate corresponding videos using the target video generation model. VisionReward then evaluates each video and assigns a reward score. This process results in $(x,p_{m},r_{m})$, where $r_{m}$ is the score from VisionReward. By comparing these scores, we determine which prompts lead to better video quality. Assuming $r_{m}<r_{m+1}$, we can then construct the video-level DPO pair, $(x,p_{m}<p_{m+1})$, resulting in $D_{\text{video}}$.
Model Training.
As we get text-level and video-level preference pairs, we can conduct DPO on the SFT model. The training dataset is denoted as $D_{\text{dpo}}=D_{\text{text}}\cup D_{\text{video}}$ and the DPO loss is described as follows:
$\mathcal{L}_{\text{DPO}}(\pi_{\theta};\pi_{\text{ref}})=$ $-\mathbb{E}_{(x,p_{w},p_{l})\sim D_{\text{dpo}}}\Bigg{[}\log\sigma\Bigg{(}\beta\log\frac{\pi_{\theta}(p_{w}|x)}{\pi_{\text{ref}}(p_{w}|x)}$ (2)
$-\beta\log\frac{\pi_{\theta}(p_{l}|x)}{\pi_{\text{ref}}(p_{l}|x)}\Bigg{)}\Bigg{]}$
where $\pi_{\theta}$ is the policy model initialized with the SFT model, and $\pi_{ref}$ is a fixed reference model also initialized with the SFT model. Here, $p_{w}$ and $p_{l}$ represent the chosen ($p_{j_{\text{refined}}}$ from $D_{\text{text}}$ and $p_{m+1}$ from $D_{\text{video}}$) and rejected ($p_{j}$ from $D_{\text{text}}$ and $p_{m}$ from $D_{\text{video}}$) prompts, respectively. This process results in the final DPO model, which is capable of generating prompts that improve video quality while maintaining safety and accuracy.
Method MonetBench VBench
Align-
ment
Stability
Preser-
vation
Physics Overall
Human
Action
Scene
Multiple
Objects
Appear.
Style
CogVideoX-2B
Original Query 1.11 0.25 0.56 0.31 3.27 80.00 28.34 40.17 22.60
Promptist 0.88 0.25 0.55 0.29 2.87 67.40 18.37 27.44 23.12
Prompt-A-Video 1.23 0.27 0.61 0.33 3.58 90.60 43.85 68.26 22.33
GLM-4 Few-Shot 1.28 0.27 0.59 0.33 3.57 96.20 55.51 68.40 23.47
GPT-4o Few-Shot 1.26 0.27 0.58 0.33 3.58 98.20 52.53 63.63 23.73
VPO-SFT 1.28 0.28 0.60 0.33 3.59 97.00 55.04 68.98 24.13
VPO w/o TL FDBK 1.32 0.29 0.62 0.33 3.72 96.40 54.78 67.79 24.15
VPO 1.34 0.29 0.63 0.34 3.76 99.00 55.83 70.17 24.20
CogVideoX-5B
Original Query 1.31 0.29 0.62 0.34 3.77 88.00 41.32 45.67 23.37
Promptist 1.08 0.28 0.62 0.33 3.42 77.40 24.93 18.34 23.27
Prompt-A-Video 1.42 0.31 0.66 0.35 4.05 91.80 45.40 74.41 22.63
GLM-4 Few-Shot 1.46 0.29 0.64 0.35 3.98 98.40 55.60 72.38 24.39
GPT-4o Few-Shot 1.48 0.29 0.64 0.34 4.03 99.20 53.13 72.21 24.20
VPO-SFT 1.47 0.30 0.65 0.35 4.01 97.20 58.40 73.70 24.55
VPO w/o TL FDBK 1.52 0.31 0.67 0.35 4.12 97.60 54.59 72.99 23.96
VPO 1.52 0.31 0.67 0.36 4.15 99.60 55.68 75.73 24.57
Table 1: Main results on MonetBench and VBench (%). ‘Appear. Style’ stands for Appearance Style. VPO w/o TL FDBK indicates that only video-level rewards are used during the preference optimization stage. The highest results for each video generation model is bolded.
Figure 3: Manual evaluation results comparing VPO with baseline methods. VPO demonstrates a significant advantage in instruction-following, video quality, and overall performance.
4 Experiments
4.1 Experimental Setup
To thoroughly evaluate the effectiveness of VPO, we conduct a series of comprehensive experiments covering multiple aspects, including the main results on text-to-video benchmarks, text-level alignment, safety evaluation, and comparisons with RLHF methods for video generation models. Additionally, we assess the iterative improvement capability, as well as the generalization performance of VPO across different video generation models. Implementation details are described in the Appendix.
Backbone Models & Baselines.
Our experiments are conducted on several popular video generation models, including the CogVideoX series and Open-Sora 1.2. The CogVideoX series excel in producing coherent, motion-rich, long-duration videos, while Open-Sora 1.2 is an efficient open-source alternative. For CogVideoX, we use the official prompt optimization method based on GLM-4 in-context learning, along with GPT-4o in-context learning as baselines. For Open-Sora 1.2, we adopt the official GPT-4o rewriting method as our baseline. Additionally, we introduce an alternative baseline using the image prompt optimizer, Promptist [10]. Moreover, we compare VPO with Prompt-A-Video [13].
Evaluation Benchmarks.
We use VBench [12] and MonetBench [27] for general text-to-video evaluation. VBench assesses multiple aspects, including Human Action, Scene, Multiple Objects, and Appearance Style. MonetBench evaluates diverse scenarios with metrics such as Alignment, Stability, Preservation, and Physics. For query alignment, we curate 500 diverse queries using LLM-based filtering and manual verification. GPT-4o is leveraged to assess whether the optimized prompts violate our predefined principles, such as containing harmful content or omitting key details. For safety, we use a subset of T2VSafetyBench [16]. More details are provided in the Appendix.
4.2 Text-to-Video Benchmark Results
As shown in Table 1, we implement VPO on CogVideoX-2B and CogVideoX-5B. In both cases, VPO significantly outperforms baseline approaches. The obvious gap between directly using the original query with using prompt optimization techniques highlights the critical role of prompt optimization in video generation. Furthermore, the difference between VPO-SFT and VPO models demonstrates the necessity of incorporating multiple feedback preference optimization. Notably, the superior performance of VPO over VPO without text-level feedback indicates that improving safety and accuracy at the text level enhances general text-to-video tasks.
To further validate the effectiveness of VPO, we conduct a pairwise human evaluation on CogVideoX-5B. We ask annotators to focus on the following aspects:
  • Instruction-Following: The degree to which the generated video adheres to the user query.
  • Video Quality: The quality of generated videos, like coherence, stability, and adherence to physics.
  • Overall Preference: A holistic judgment based on instruction-following and video quality.
The results are shown in Figure 3, which are consistent with evaluation on benchmarks. VPO significantly surpasses other methods. The superior performance of VPO over VPO without text-level feedback on the alignment dimension indicates that a more accurate prompt optimizer could help increase the instruction-following abilities of video generation models.
Method Aligned $\uparrow$ Misaligned $\downarrow$
Unsafe Imprecise Refusal
GLM-4 Few-Shot 83.4 5.4 10.0 1.2
GPT-4o Few-Shot 86.4 2.4 8.6 2.6
VPO-SFT 93.8 0.8 5.4 0.0
CogVideoX-2B
VPO w/o TL FDBK 93.0 2.0 5.0 0.0
VPO 94.6 0.6 4.8 0.0
CogVideoX-5B
VPO w/o TL FDBK 92.8 1.2 6.0 0.0
VPO 94.8 0.4 4.8 0.0
Table 2: Evaluation results (%) of query alignment. We highlight the best results for each video generation model in bold.
Figure 4: Manual evaluation results on safety tasks. VPO demonstrates substantially improved harmlessness, particularly in terms of the completely safe rate (Level 1).
4.3 Text-Level Alignment
We further evaluate VPO on a more complex test set constructed from real-world user queries, focusing on accuracy and harmlessness. As shown in Table 2, VPO largely enhances text-level alignment, surpassing alternative methods, including much larger LLMs. Importantly, our findings reveal a critical limitation of relying solely on LLM-based in-context learning for prompt optimization: these models may refuse to process queries containing sensitive keywords or abstract contents, such as ”20 - 11 coins = 9 coins.” This underscores the necessity of training a comprehensive prompt optimizer capable of dealing with diverse user queries.
4.4 Safety Evaluation
We conduct a manual safety evaluation using a subset of T2VSafetyBench, where annotators assign safety scores to generated videos on a four-point scale:
  • Level 1: Completely safe.
  • Level 2: Safe but contains minor unsafe elements.
  • Level 3: Unsafe.
  • Level 4: Extremely unsafe.
As shown in Figure 4, VPO substantially improves video safety compared to LLM few-shot methods. This enhancement is particularly evident at Level 1, where VPO ensures a significantly higher proportion of completely safe outputs. Importantly, incorporating text-level feedback further enhances safety, whereas neglecting this component leads to a lower safety rating than the SFT model, indicating that solely optimizing for video quality could damage safety.
Figure 5: Pairwise evaluation results using VisionReward. Standard refers to the original video generation model with the official prompt rewriting method. VPO outperforms Diffusion DPO and can be combined with it to provide additional benefits.
4.5 Comparison to RLHF
We reveal that the prompt optimization method is not only important to bridge the training and inference gap of video generation models, but it is also a good way to align the video generation model with human preference. In this experiment, we compare VPO with Diffusion DPO [25]. Diffusion DPO is applied with the official GLM-4 prompt rewriting. Using VisionReward for pairwise evaluation, we find that both methods enhance CogVideoX-5B’s performance, with VPO outperforming Diffusion DPO, as shown in Figure 5. Notably, VPO and Diffusion DPO can be combined for additional gains, demonstrating that optimizing both the video generation model and prompt optimizer can lead to superior results.
Figure 6: (Left) Case study on a harmful query; (Right) Case study on a harmless query. Some prompts are omitted due to space constraints.
4.6 Generalization across Models
As the preference optimization stage of VPO relies on a specific video generation model to construct DPO pairs, we would like to investigate if the prompt optimizer could generalize to other models. To test this, we apply the prompt optimizer trained on CogVideoX-2B to Open-Sora 1.2. As shown in Table 3 and Table 4, we demonstrate that VPO could generalize across different video generation models, yielding non-trivial performance improvements on both VBench and MonetBench.
Method
Human
Action
Scene
Multiple
Objects
Appear.
Style
Original Query 88.80 44.08 55.99 23.87
GPT-4o Few-Shot 92.40 53.21 65.02 23.84
VPO-SFT 95.80 51.41 64.28 23.86
VPO 97.00 53.58 67.88 23.93
Table 3: Evaluation results (%) of Open-Sora 1.2 on VBench.
Method
Align-
ment
Stability
Preser-
vation
Physics Overall
Original Query 1.01 0.21 0.54 0.31 2.90
GPT-4o Few-Shot 1.12 0.21 0.55 0.31 3.07
VPO-SFT 1.12 0.23 0.56 0.31 3.08
VPO 1.13 0.24 0.58 0.33 3.18
Table 4: Evaluation results of Open-Sora 1.2 on MonetBench.
5 Case Study
Figure 6 presents a comparative analysis of VPO and the GLM-4 Few-Shot approach in handling both harmful and harmless queries. On the left, the original query describes a disturbing hospital scene with exposed wounds, blood, and medical waste. The few-shot method generates a detailed description of these unsettling elements, leading to a harmful video. In contrast, VPO shifts the focus to a professional and sterile hospital environment, emphasizing care and resilience rather than distressing imagery. This demonstrates VPO’s ability to refine harmful prompts into ethical and appropriate content while maintaining contextual relevance. On the right, we examine a harmless query about hugging. While the few-shot approach generates a semantically rich scene, the resulting video suffers from poor quality, with unnatural body movements and misplaced limbs. In comparison, VPO produces a more natural and visually coherent video, with smoother and more realistic movements. This highlights the importance of integrating feedback from generated videos to build a helpful prompt optimizer. These cases show that VPO not only improves video quality but also ensures that the generated content remains safe and appropriate. More examples are provided in the Appendix.
6 Conclusion
In this work, we introduce VPO, a principle-driven framework for optimizing prompts in video generation, ensuring harmlessness, accuracy, and helpfulness. Unlike conventional in-context learning approaches, which rewrite prompts without considering their actual effectiveness, VPO improves video generation quality through principle-based SFT and multi-feedback preference optimization, incorporating both text-level and video-level feedback. Through extensive experiments, we demonstrate the superiority of VPO over the baseline methods, delivering safer, more accurate, and higher-quality video generation, effectively bridging the gap between training and inference in video generation models. Moreover, we reveal that VPO also serves as an effective RLHF method and can be integrated with RLHF methods on video models, bringing orthogonal improvements. Additionally, we demonstrate the generalization of VPO across video models, showing the potential for developing a general text-to-video prompt optimization model.
7 Acknowledgement
This work was supported by the National Science Foundation for Distinguished Young Scholars (with No. 62125604). This work was also supported by Tsinghua University Initiative Scientific Research Program. We would also like to thank Zhipu AI for sponsoring GPU computing and API cost consumed in this study.
Appendix
8 Prompt Template for Data Construction
We show the prompt template for constructing the principle-based SFT dataset in Figure 7 and Figure 8. The prompt template shown in Figure 8 is also used for constructing preference pairs.
Figure 7: Prompt template for the initial optimized prompt few-shot generation.
Figure 8: Prompt template for critique and refinement generation.
9 Implementation Details.
In our experiments, we use LLaMA3-8B-Instruct [7] as the base model to train the prompt optimizer. Both the SFT and DPO stages utilize approximately 10k queries for data construction, including around 1k safety-related queries. For SFT data construction, GPT-4o is used to generate optimized prompts, provide critiques, and refine the optimized prompts. Detailed prompts are provided in Section §8. For SFT training, we set the learning rate to 2e-6 and train for five epochs. The training employs a 0.1 warmup ratio and a batch size of 64. The AdamW optimizer is used with $\beta_{1}=0.9$ and $\beta_{2}=0.999$. In the DPO stage, we sample 4 prompts for each query with a temperature of 0.9. For text-level preference pairs, we also choose GPT-4o to judge and refine. In the DPO stage, we sample four prompts per query with a temperature of 0.9. For text-level preference pairs, GPT-4o is used for judgment and refinement. For video-level preference pairs, we ensure they adhere to text-level principles, selecting prompt pairs with a reward score difference greater than 0.5. This process generates approximately 5k preference pairs for DPO training. The DPO training is performed with a learning rate of 5e-7, $\beta=0.1$, a 0.1 warmup ratio, and a batch size of 16 for one epoch. For both SFT and DPO training, we utilize the DeepSpeed Zero-3 strategy [19]. All experiments are conducted on $\times8$0G NVIDIA H800 GPUs.
MonetBench comprises seven content categories and thirteen challenge categories, covering a broad range of video scenarios and creative aspects. For the evaluation of query alignment, we employ both automatic filtering via GPT-4o and manual verification to ensure a diverse set of 500 queries spanning varying difficulty levels. The T2VSafetyBench assesses safety risks across 12 dimensions. Since GPT-4o often refuses to provide judgments, we manually evaluate a subset of 200 samples while preserving the original data distribution.
10 Comparison with VBench Long Prompts
We show the comparison with VBench Long Prompts in Table 5. VPO consistently outperforms the VBench Long Prompts baseline.
Method
Human
Action
Scene
Multiple
Objects
Appear.
Style
VBench Long Prompts (2B) 98.00 51.33 63.81 24.07
VPO (2B) 99.00 55.83 70.17 24.20
VBench Long Prompts (5B) 98.40 53.67 65.67 24.41
VPO (5B) 99.60 55.68 75.73 24.57
Table 5: Comparison with VBench Long Prompts.
Method
Align-
ment
Stability
Preser-
vation
Physics Overall
Original Query 1.31 0.29 0.62 0.34 3.77
Iteration 1 1.52 0.31 0.67 0.36 4.15
Iteration 2 1.53 0.31 0.67 0.35 4.17
Iteration 3 1.52 0.32 0.68 0.36 4.18
Iteration 4 1.51 0.31 0.67 0.37 4.17
Table 6: Evaluation results of iterative improvement of VPO on MonetBench.
11 Iterative Improvement
As VPO could optimize user input for better results, a natural problem arises: can we iteratively improve the prompt for higher-quality videos? The answer is yes. We iteratively optimize the user’s query for four iterations and find that the performance improves in the first three iterations and then becomes stable, as shown in Table 6. This also shows an important characteristic: the prompt optimizer will not damage the performance in further optimization. It would like to preserve the prompt if it is already good enough.
12 Case Study
In this section, we present case studies of VPO compared to other baseline methods. Figure 9 shows a scenario involving a harmless query: ”A person is cheerleading.” The original query, which is short and simple, poses a challenge for video generation models, making it difficult to produce high-quality results. While few-shot methods generate more detailed captions, they still fail to produce stable and high-quality videos. This highlights the importance of considering the final video quality when optimizing user queries. In contrast, VPO consistently generates stable and visually appealing videos, outperforming other methods. Figure 10 depicts a harmful query involving a scene where a person falls onto the tracks, staining them red. The baseline methods generate unsafe content, such as blood on the tracks, emphasizing the need for safety alignment during the prompt optimization process. Notably, VPO without text-level feedback (denoted VPO w/o TL FDBK) also produces a harmful scenario, where the train runs over the person, and the tracks turn red. This emphasizes the crucial role of text-level feedback in the preference optimization process.
Figure 9: Case study of the performance of VPO.
Figure 10: Case study of the performance of VPO on safety task.